OpenStudy (calculusxy):

When starting a new unit, a teacher announces that each day after the first, the class must do twice the total number of problems that had been assigned on all previous days. The class works for 6 days and on the 7th day she says that this is our last day of the unit. On that last day, what fraction of the problems do they still have to complete. @hartnn

OpenStudy (anonymous):

Would you agree that the number of problems increases like this: 1, 2, 6, 18, 54, 162, 486

OpenStudy (calculusxy):

What is the rule that is in this sequence?

OpenStudy (anonymous):

2 times the sum of all numbers of questions previously assigned

OpenStudy (anonymous):

\[a_n = 2\sum_{i=1}^{n-1}a_i\]

OpenStudy (calculusxy):

What is this formula? I have no idea what this means.

OpenStudy (anonymous):

suppose the teacher assigns 1 problem on the first day. how many will the teacher assign on day 2?

OpenStudy (calculusxy):

2?

OpenStudy (anonymous):

okay, so now how many problems have been assigned?

OpenStudy (anonymous):

in total.

OpenStudy (calculusxy):

If you are talking about how many problems have been answered in Day 2, then I am going with the total of 2 problems.

OpenStudy (anonymous):

no, from the beginning, just like it says in the problem.

OpenStudy (calculusxy):

Is the sequence going to be like 1,2,4,16... (Somewhere along with that rule?)

OpenStudy (anonymous):

please, just stick to the question i asked... i'm trying to help you see the pattern

OpenStudy (calculusxy):

I really don't understand your question, I don't get by what you mean by" okay, so now how many problems have been assigned?"

OpenStudy (anonymous):

1 question assigned the first day, yes? 2 questions assigned the second day, yes? since the beginning, how many questions have been assigned?

OpenStudy (calculusxy):

But we don't know if they really assigned 1 question on the first day. Wouldn't that be like a variable?

OpenStudy (anonymous):

doesn't matter, we can multiply by n

OpenStudy (anonymous):

n, 2n, etc.

OpenStudy (calculusxy):

Ok would that be like n, 2n, 4n, 16n,and so on? Then can solve for the variable.

OpenStudy (anonymous):

no... again: please just answer my question. i'm trying to help you see the pattern. you obviously have the wrong pattern stuck in your head. i suggest you remove that, open your mind and please answer my question.

hartnn (hartnn):

the pattern 1,2,4,8,16,.... is correct, right ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

1, 2, 6, 18, 54, 162, 486

hartnn (hartnn):

" twice the total number of problems that had been assigned on all previous days" i am sorry, you're right

OpenStudy (anonymous):

no worries!

OpenStudy (anonymous):

or, n, 2n, 6n, 18n, 54n, 162n, 486n, ...

hartnn (hartnn):

yeah, even with n=1, you'll get the same final answer, so just go with 1,2,6,18...

OpenStudy (calculusxy):

Can you give me a relevant formula for this pattern?

OpenStudy (anonymous):

also since \(a_n=2\sum_{i=1}^{n-1}a_i\) and the total number of problems is \(a_n+\sum_{i=1}^{n-1}a_i\) we can conclude that proportion of problems remaining \[=\frac{a_n}{a_n+\sum_{i=1}^{n-1}a_i}=\frac{2\sum_{i=1}^{n-1}a_i}{2\sum_{i=1}^{n-1}a_i+\sum_{i=1}^{n-1}a_i}=\frac{2\sum_{i=1}^{n-1}a_i}{3\sum_{i=1}^{n-1}a_i}=\frac{2}{3}\]

OpenStudy (calculusxy):

Can you please translate this? I am not up to that level yet.

OpenStudy (anonymous):

can't seem to find a clever way to give a formula for this one. it's nested sums and i'm a bit rusty on all of this.

OpenStudy (anonymous):

\[a_1 = 1 \Rightarrow a_2 = 2\sum_{i=1}^{1}a_i = 2(a_1) = 2(1) = 2\] \[\Rightarrow a_3 = 2\sum_{i=1}^{2}a_i = 2(a_1+a_2) = 2(a_1+2a_1) = 2(1+2)=6\] \[\Rightarrow a_4 = 2\sum_{i=1}^{3}a_i = 2(a_1+a_2+a_3) = 2(a_1+2a_1+2(a_1+2a_1)) = 2(1+2+6)=18\] etc...

OpenStudy (calculusxy):

What do you mean. By (2a-1)...

OpenStudy (anonymous):

not \(2a-1\text{, but }2\cdot a_1\)

OpenStudy (calculusxy):

Yeah so what does that mean?

OpenStudy (anonymous):

2 times \(a_1\). remember, \(a_1 = 1\)

OpenStudy (calculusxy):

What do all those numbers by the big E mean?

OpenStudy (anonymous):

don't worry about it. what level is your class?

OpenStudy (calculusxy):

Um seventh

OpenStudy (calculusxy):

Ur's?

OpenStudy (anonymous):

okay. you won't be seeing this for a while. have a look if you're really interested... http://www.purplemath.com/modules/series.htm

OpenStudy (anonymous):

i have a degree in math

OpenStudy (calculusxy):

BA or MA?

OpenStudy (anonymous):

BA but did about 35 units towards an MA

OpenStudy (calculusxy):

Almost with MA?

OpenStudy (anonymous):

yeah

OpenStudy (calculusxy):

Thanks for ur help. But how did u get 2/3?

OpenStudy (anonymous):

do you understand how to get the sequence? 1, 2, 6, 18, 54, 162, 486

OpenStudy (anonymous):

if so, the total number of problems is 1+2+6+18+54+162+486 = 729 486/729 = 2/3